# Phase segregation lubricates paste extrusion

### v1

@@ Abstract
Extrusion is a critical process in the manufacturing of various organic materials, including animal feed [1], human food [2], pharmaceuticals [3], and biomass for bioenergy [4]. During extrusion, a dense particle-fluid mixture, or paste, is pushed through a narrow channel under high pressure [5, 6]. Despite its widespread use, many questions remain about how particle-fluid interactions and frictional forces at the wall influence high-pressure paste flow [7, 8], particularly whether laboratory-scale findings translate to more complex industrial settings[9]. Here, we quantify for the first time how a micrometer thin fluid lubrication layer emerges from particle-fluid interactions during the extrusion of organic biomass, controlling friction at the paste-channel interface. This lubrication layer forms through a pressure-induced phase segregation process [10, 9, 11] and is significantly thinner than the particle size [12], affecting their contact dynamics, making friction at the interface tunable. Our results reveal how the thickness of the lubrication layer is influenced by paste composition, enhancing our fundamental understanding of phase segregation during industrial extrusion processes and offering potential pathways for reducing energy consumption and product failure  [13] in the industrial material production.
@@
The flow behavior of concentrated suspensions and pastes remains a challenge to understand, affecting processes ranging from human food production and additive manufacturing to medical and clinical applications [10, 14, 15, 16, 17, 18, 11, 19, 2]. Pastes are mixtures of particles and a small amount of liquid. During their extrusion, these two-phase mixtures of particles and liquid are pressured into a pre-shaped channel, the die, to obtain a final product. Pressures used are 100s of MPa, high enough to induce particle deformation and bonding [20]. Optimization of the material and factory performance during this process happens via paste ingredient adjustments, die shape modification and process temperature and pressure alteration  [21, 6] and is performed at a daily basis in many plants worldwide. Our fundamental microscopic understanding of the typically non-Newtonian flow behavior of pastes [7, 8] during high pressure extrusion is however very limited. Controlling the extrusion process is challenging due to the numerous tuning parameters involved [5, 6], yet currently there is no physico-chemical framework that allows us to understand their interrelatedness. Here we provide one new fundamental paste-flow mechanism that binds many extrusion parameters together. We show that many extrusion tuning parameters are fundamentally linked in their ability to tune an ultrathin liquid film that emerges during extrusion. The film is only micrometers thick, as it is located between the extrusion die and the extrudate. The frictional role of the film so determines the mechanical energy usage and temperature evolution during the extrusion process, thereby also controlling a large number of product quality parameters.
@@
During paste extrusion, physical forces act on the solid-liquid mixture, deforming the particles and causing the solid particles and liquid fraction to undergo relative displacement. These phenomena often referred to as “migration”, “demixing” or solid-liquid phase separation. In general, such solid-liquid separation is known to result in flow blockages, product inhomogeneity and product failure [17, 22, 23, 13]. It was conjectured long ago that also a beneficial liquid-rich layer can form between the wall of the die channel and the flowing paste [10], resulting in self-lubricating paste flow. Such liquid lubrication layers are known to change sliding resistance by orders of magnitude [24, 25], particularly in the so-called “mixed”-regime where a lubrication layer starts to reduce the solid-to-solid sliding friction [26]. As the lubrication layer affects friction or paste tribology, it influences the mechanical energy use and heat generated inside the extrusion channel at the (sub)micrometer length scale. In the extrusion die, precise tuning of both energy and temperature are necessary to induce and control agglomeration of the particulate mixture through various physical and chemical transitions, such as chemical reactions, melt transitions, and sintering [20]. Control over these transitions is crucial in the formation of a high-quality end product [20, 6, 1, 27, 2]. The control over phase segregation induced lubrication is thus an essential step in paste extrusion. Our post-extrusion surface liquid evaporation measurements quantify this phase segregation by measuring the lubrication film thickness, allowing for deep insight into the flow dynamics of pastes under pressure.
@@
Figure 1: Lubrication layer thickness increases with water use during extrusion of biomass pellets. a, During pellet manufacturing, ingredient mixtures are extruded through holes within a metal ring-die, forming cylindrically shaped pellets. The water from the steam conditioning ends up on the surface of the pellets in the high pressure extrustion due to solid-liquid phase segregation. The thickness of the lubrication layer is determined immediately after extrusion through a controlled evaporation measurement on freshly sampled pellets, recording sample mass and relative air humidity ($\mathrm{RH}$). b, Evaporation measurements show that an initial saturation period ($\mathrm{RH}>95$%, denoted as $t_{sat}$) is present, arising from the evaporation of water exposed at the pellet surface. c, The mass-detected average amount of evaporated water depends on the amount of steam added during the conditioning step: the change in water content ($\Delta\mathrm{mc}$) of $0.057$ (top), $0.033$ (middle), and $0.009$ (bottom), given in kg water per kg ingredients. The bold lines highlight the amount of water that evaporates during $t_{\mathrm{sat}}$ as indicated in (b) and hence quantifies the amount of surface water that evaporates from the pellets exterior surfaces. d, The total surface water evaporation indicates the lubrication layer thickness. The layer thickness is micrometers thick and increases as more water is added in the steam conditioning process as captured by $\Delta\mathrm{mc}$. e, Surface topology scans of dried pellets confirm thickness calculations, as they do not show striations along the extrusion direction when the lubrication layer is thicker than the intrinsic height of the asperities in the metal die (top). Decrease of the lubrication layer thickness results in clear striations on the pellets surface along the extrusion direction—running left to right (bottom). Scale bar: 50 $\mathrm{\mu m}$. These images confirm that the metal die and the pellet are largely separated by the lubrication layer, schematically represented in (f).
@@
The lubrication layer thickness was suggested to be only micrometers thick [10, 7, 8, 2], a layer too thin to measure directly in any realistic extrusion process. It has thus so far been impossible to understand how this lubrication layer emerges or depends on ingredients and process parameters. Various experimental methods have been developed to analyze phase segregation in extrusion processes at laboratory scale [10, 5, 28, 29, 30, 31, 32]. These studies revealed that flow behaviour during extrusion is highly dependent on ingredient and process variables, including solid-liquid mixing ratio, particle size, deformability of particles, flow rate, pressure, and viscosity [15, 5, 33], yet stopped short of directly showing the existence of such a lubrication layer, let alone quantify it. Our surface liquid evaporation measurements [34], directly reveals (1) that this lubrication layer exists, (2) that it is indeed only several micrometers thick, and (3) we provide evidence that the thickness is set by process parameters. We (4) confirm that lubrication tuning indeed results in frictional changes by observing energy savings and temperature modulation. We demonstrate our results via pilot-scale pelleting plant trials, with realistic organic ingredient streams as typically used in biofuel and animal feed production. Our results can be translated to similar unit operations used in ceramics, foods, and pharmaceuticals production.
I Water induced lubrication in pelleting processes
@@
Pellets are formed in a unit operation known as the pellet extrusion process (Fig. 1a&f), where particle mixtures, often from multiple ingredients, are pressed by a roller through cylindrically shaped holes in a metal ring-die. The pellet extrusion process transforms the loose and low-density particle mixture into densely packed solid pellets, which are subsequently cooled and dispatched to the end-user. Unlike hard inorganic particle pastes used in e.g. concrete extrusion, organic plant-based particles utilized in food and feedstuff production exhibit hygroscopic properties, are capable of binding or releasing water, and undergo deformation and physico-chemical transitions while being processed. Consequently, the interaction mechanism and physical forces at the boundary layer between the metal extrusion die and the food/feed paste [2] are expected to differ from those of hard particles which for example are subject to excluded volume effects at the die-wall[7, 5, 8, 35].
@@
Prior to pelletization, conditioning uniformly mixes the ingredient particles and steam. The steam condenses upon cooling, transferring its latent heat into the particle mix and forms water [1, 36, 37]. Consequently, the steam conditioning process sets the temperature and amount of water present in the paste before it enters the extrusion die. It is known that increasing the amount of water in the mixture reduces the mechanical energy required per kilogram of produced pellets [37]. While the water-induced lubrication in pellet extrusion is evident, the underlying mechanism remains unclear. The observed layer is too thin to be related to a particle migration phenomena, although such effects can also significantly change effective paste viscosity [38, 39]. Hitherto, one main assumption is that the paste during pellet extrusion internally retains its uniform moisture distribution from right after the mixing/conditioning step [40, 41]. We show that this is not the case: the lubricating effects of added water is caused by the formation of a thin lubricating film that partially fills the pores between the metal die and the sliding pellet (Fig. 1f), reducing the real contact area between those respective surfaces, as expected in the mixed-lubrication regime [42]. Our perspective also explains the observed reduction in equipment wear with added water [37].
Figure 2: Experimental and numerical drying results support the hypothesis of a lubrication layer. Two types of pellets underwent drying using our model drying setup (solid lines), complemented by results obtained through numerical modeling (dashed lines). The results are presented as the pellet moisture ratio, the normalized mass decay (unitless) over $900$ seconds of drying (see Eq. 6). In yellow: results corresponding to the experimental drying of pellets with a uniform moisture content, achieved through vapor sorption (stored for 1 week at over $95\%$ RH), and numerical simulation considering an initially uniform moisture distribution. In blue: results corresponding to the experimental drying of fresh pellets immediately after extrusion and numerical simulation for a pellet with a high moisture content localized at its surface, as schematically depicted. The initial drying rate is strongly affected by the presence of surface water, resulting in a quick drying profile, in contrast to pellets with a uniform moisture distribution, for which the rate of drying is determined by the transport by diffusion. Additional results are shown in Fig. A.1.
II Formation of lubrication layers during paste extrusion
@@
We provide evidence for the existence of the lubrication layer through a pilot-scale pelleting trial. Approximately 18 tons of pellets were produced while systematically adjusting the conditioning settings, as described in the Methods section. The fresh pellets were sampled immediately after extrusion and placed in a custom designed sample container permitting drying of the product (Fig. 1a). The drying method involves exposing a sample container of pellets to a continuous, known flux of fresh air and recording the inlet and outlet air temperature and relative humidity (RH), along with the sample mass [34]. In 14 out of 15 drying experiments, a saturation period was observed ($t_{\mathrm{sat}}$), during which the passing air reached saturation (Fig. 1b). Over this period, up to 40% of the water introduced by conditioning was lost through evaporation (Fig. 1c), corresponding to the moisture “flash-off” commonly observed in pelleting factories [1]. Our modeling demonstrates that the saturation of the drying air and the initial rate of drying, observed during the drying of freshly produced pellets, can only occur if the evaporating water comes directly from the pellet surface. This is in contrast to pellets with an elevated but uniform moisture profile, which can be generated through vapor sorption (see Fig. 2; we refer to Fig. A.1 b&c and SI Section: A.2 for a detailed description of the numerical model and the results).
@@
Evaporation of the lubrication layer thus causes the initial saturation of drying air (indicated in Fig. 1b). The amount of water evaporating during this period can be quantified as we know the total surface area of the pellets in the drying experiment from the drying mass, density and shape. We so obtain the average water layer thickness, $\lambda$, on the pellets surfaces (Fig. 1d). The thickness of the lubrication layer increases from a few micrometers, similar to the height of surface asperities on the metal die (Fig. 1e) in which case we expect frictional interactions that are also known as Stribeck boundary lubrication [24]. As we increase water content of the paste ($\Delta\mathrm{mc}$) by increasing the steam flow in the conditioning process, the layer thickness increases to several tens of micrometers. Compared to the smallest individual particles produced by the milling procedure, this layer remains very thin [12] (Fig. 1d). The water layer thus only partially separates the pellet from the die, reducing the area of direct contact between the pellet and die surface and so leads to less damage to the particles embedded in the pellet surface (Fig. 1e&f and SI section: A.4) and to less wear in the die. As the lubrication layer grows, we so expect to enter the “mixed regime” lubrication in the Stribeck perspective, and a concomitant reduction in friction. The presence of the lubrication layer also introduces plug flow in the compressed biomass as there is no visible internal deformation, uniform density and negligible porosity (see supportive X-ray tomography evidence described in SI Section: A.3). These collective findings suggest that the lubrication layer in organic feedstuffs may not arise from factors like particle size distribution and the excluded volume effect, as observed in hard particle pastes. Instead, it appears to be influenced by the roughness of the die wall. In this scenario, water, with a much lower viscosity compared to the feedstuff, flows into the die’s cavities and reduces the shearing of the feedstuff at the paste-die-wall interface by occupying part of the cavity volume, thus making it inaccessible for the feed particles. Consequently, the real contact area between the paste and metal die reduces, and the energy use decreases, as shearing of the water layer is much more favorable than shearing of the feedstuff material along the rough metal wall. This hypothesis aligns with industrial findings indicating that new pellet dies require more energy during pellet extrusion than used dies [43]. We attribute this energy use reduction over time to a decrease in the paste-die-wall contact area as the die’s largest asperities smooth out through wear. By smoothing of the surface, the water film can cover a larger area (the ratio of $\lambda$ over the wall roughness increases), further reducing the real contact area between the paste and die [25].
@@
The ability to measure and monitor the lubrication layer thickness enables us to decouple process dependent parameters such as steam use, die-hole dimensions and production rate (the latter, together with die-geometry specifications, determines residence time) from ingredient mixture dependent parameters, such as its chemical composition, particle size, particle size distribution and thermomechanical properties [44]. We identify two important consequences: we can now identify the source of mechanical energy use of pellet extrusion—the friction at the paste-die-wall interface—and consequently reduce it. Second, by tuning the lubrication layer, we can adjust the amount of heat added during extrusion, giving control over product treatment and quality.
III The tribological role of the lubrication layer during pellet extrusion
During pellet extrusion, the sliding friction between the pellet and the metal die contributes to the energy use of the pellet press. This energy use is proportional to the extrusion pressure and increases exponentially with the compression ratio—the die-length over the die-diameter ratio—of the extrusion die $L/D$ and with the apparent coefficient of friction between the die and the paste, $\mu$ [45, 46, 47] (see SI section A.4). The dry solids friction is determined by the paste composition and its interaction with the die’s metal surface [48, 49]. In the case of animal feed, this composition is set by the raw ingredient mixture. This dry coefficient of friction can not be actively manipulated during pellet extrusion, yet the formation of a thin lubrication layer can contribute to a smooth transition from the “boundary lubrication regime” into the so-called “mixed regime” [25], significantly reducing the coefficient of friction (see SI section A.4). The decrease in the apparent coefficient of friction, $\mu$, has a direct impact on the specific mechanical energy use of the pellet press—the amount of mechanical energy required for the production of a kilogram of pellets (Fig. 3b). Hence, adjusting the lubrication layer through steam addition enables modification of the friction coefficient within a running production line, without altering the ingredient mixture composition or press/die configuration. As a consequence of the lubrication, frictional heat changes the temperature gain in pellets over the extrusion die, here called $\Delta T_{\rm die}$. We show this effect in Fig. 3c, which affects final product properties (such as “durability” as detailed in SI section A.5). An increase in extrusion rate also suppresses friction and hence SME, in accordance to Stribeck expectations [25] (see SI Fig. A.8) Temperature fluctuations can change the physico-chemical properties of the organic ingredient streams, which may result in gelatinization and melting of starches and denaturation of proteins, which subsequently affect the nutritional quality and availability during consumption [36]. Measurement and subsequently tuning of the lubrication layer is thus essential for further optimization of pelleting processes as it enables a more precise control over the mechanical energy use and product temperature compared to the current set of operating variables.

### v2

@@ Abstract
Dense particle-fluid mixtures, or pastes, are encountered in the production of various materials, including animal feed [1], human food [2], pharmaceuticals [3], and biomass for bioenergy [4]. The flow behavior of such dense deformable particulate media is poorly understood, as the interplay between applied stresses, particle deformability and interstitial fluids can be very complex. One challenging context is high pressure pipe flow, encountered in extrusion [5, 6]. Despite its widespread use, many questions remain about how during high pressure flow of the paste, the particle-fluid mixture behaves and interacts with boundaries [7, 8, 9]. We show how high pressure paste extrusion induces the formation of a fluid boundary thinner than the particle size [10]. The induced fluid layer emerges from a pressure-induced phase segregation process [11, 9, 12]. The fluid layer is sufficiently thin to affect particle-wall contacts, making the paste friction coefficient tunable. Our results so offer potential pathways for reducing energy consumption and even extrusion product composition and failure [13].
@@
Introduction
The flow behavior of concentrated suspensions and pastes plays a role in processes ranging from human food production and additive manufacturing of plastics and concrete structures to medical and clinical applications [11, 14, 15, 16, 17, 18, 12, 19, 2, 20]. During the extrusion of these mixtures of particles and a small amount of liquid, the two-phase mixture is pressured into a pre-shaped channel, usually called the die. High extrusion pressures are required to promote particle bonding in the extrudate [21] and produce a rigid end product. Despite the widespread use of pastes, the flow behavior of these materials still remains poorly understood. Due to the pressures involved, the flow of dense pastes is often characterized as plug flow [7, 8], referring to the perspective that the dense particle phase moves through the die as a rigid body. Yet, these high-pressures conditions in extrusion complicate paste flow substantially: the high pressure flow environments are challenging to probe experimentally, but the elevated stress levels also fundamentally changes the paste flow dynamics.
In particular, for extrusion context, applied pressures can be high enough to induce particle deformations. The particle deformations make the pore space between the particles change substantially over time and space during the flow. The liquid is present in the pore space, but its role in determining the overall flow dynamics of the paste is poorly understood, as most suspension physics focuses on dynamics in the particle phase [20]. High pressure paste flows are hence fundamentally different from dense non-deforming particle suspension flows. This makes our microscopic understanding of the typically non-Newtonian flow behavior of pastes very limited [7, 8]. Another challenge is the numerous tuning parameters involved [5, 6]; lacking a framework of understanding paste flows also limits our ability to understand the interrelatedness of these parameters.
Figure 1: (a) Sketch of the high pressure extrusion flow dynamics. A steam-wetted powder is pressed into a die channel. During the transport of the paste through the channel, the paste is consolidated and its liquid content migrates towards the boundary between the consolidated plug and the die wall, affecting local friction and heat exchange and final agglomerate behavior. Experimental control parameters are channel length $L$, diameter $D$ and flux $Q$. We quantify the lubrication layer thickness $\lambda$ present on the extrudate via its drying process. (b) During channel flow, the density profile of the solid $\phi_{\rm solid}$ and the liquid $\phi_{\rm liquid}$ change from homogeneous to highly depleted/concentrated at the wall ($\phi_{\rm solid},\phi_{\rm liquid}$ respectively) while the overall compressibility of ingredients is limited ($\phi_{\rm solid}+\phi_{\rm liquid}=1$).
Here we show that paste extrusion flow dynamics is associated with the formation of a liquid film on the surface of the paste. The film is thinner than the particle size [10] and emerges from the paste pore space in a phase segregation process. We indicate the main phenomenology in Fig. 1a, along with our experimental control parameters. The liquid film is only micrometers thick, and located between the extrusion die and the extrudate. The frictional role of the film determines the mechanical energy usage and temperature evolution during the extrusion process, thereby also controlling the binding process between the particles and hence a large number of product quality parameters, which we describe elsewhere [22, 23]. Optimization of the material and factory performance of extrusion processes of course happens already on a daily basis via paste ingredient adjustments, die shape modification and process temperature and pressure alteration [24, 6]. Our results allow to quantitatively connect these parameters via knowledge of the dynamics of the lubrication layer, and can so have major implications for many industrial processes.
Phase segregation
@@ Introduction
During paste extrusion, physical forces act on the solid-liquid mixture, deforming the particles and causing the solid particles and liquid fraction to undergo relative displacement. These phenomena often referred to as “migration”, “demixing” or solid-liquid phase separation. The phase migration effect is schematically indicated in Fig. 1b, highlighting how the relative local fractions of solid and liquid $\phi_{\rm solid},\phi_{\rm liquid}$ change from uniform at the channel entrance to highly localized ($\phi_{\rm liquid}$) at the end of the channel. In general, such solid-liquid separation is known to result in flow blockages, product inhomogeneity and product failure [17, 25, 26, 13]. It was conjectured long ago that there is also a beneficial effect: the liquid-rich layer can form between the wall of the die channel and the flowing paste [11], resulting in self-lubricating paste flow. Such liquid lubrication layers between rigid surfaces are known to change the surface sliding resistance by orders of magnitude [27, 28], particularly in the so-called “mixed”-regime where a lubrication layer starts to reduce the solid-to-solid sliding friction [29]. As the lubrication layer affects friction or paste tribology, the layer influences the mechanical energy use and heat generated inside the extrusion channel at the (sub)micrometer length scale. In the extrusion die, precise tuning of both energy and temperature are necessary to induce and control agglomeration of the particulate mixture through various physical and chemical transitions, such as chemical reactions, melt transitions, and sintering [21]. Control over these transitions is crucial in the formation of a high-quality end product [21, 6, 1, 30, 2]. The control over liquid surface migration induced lubrication is thus an essential step in paste extrusion. Our post-extrusion surface liquid evaporation measurements quantify this phase migration by measuring the lubrication film thickness, allowing for deep insight into the flow dynamics of pastes under pressure.
@@
Figure 2: (a) Cylindrically shaped pellets extruded through holes within a metal ring-die. The thickness of the lubrication layer is determined immediately after extrusion through a controlled evaporation measurement on freshly sampled pellets, recording sample mass and relative air humidity ($\mathrm{RH}$). (b) Evaporation measurements show an initial saturation period ($\mathrm{RH}>95$%, denoted as $t_{sat}$) due to evaporation of surface water. (c) The mass-detected average amount of evaporated water depends on the amount of steam added during the conditioning step: the change in water content ($\Delta\mathrm{mc}$) of 0.0570.0570.057 (top), 0.0330.0330.033 (middle), and 0.0090.0090.009 (bottom) in kg water per kg ingredients. Bold lines: the amount of water that evaporates during $t_{\mathrm{sat}}$ as indicated in (b), quantifying the amount of surface water that evaporates. (d) Total surface water evaporation converted to layer thickness $\lambda$, which increases as more water is added in the steam conditioning process ($\Delta\mathrm{mc}$). (e) Surface topology scans of dried pellets do not show striations along the extrusion direction when the lubrication layer is thicker than the intrinsic height of the asperities in the metal die (top). A thinner lubrication layer thickness results in clear striations on the pellets surface along the extrusion direction—running left to right (bottom). Scale bar: 50 $\mathrm{\mu m}$. These images confirm that the metal die and the pellet are largely separated by the lubrication layer, schematically represented in (f).
@@
The lubrication layer thickness was suggested to be a layer too thin to measure directly in any realistic extrusion process [11, 7, 8, 2]. It has thus so far been impossible to verify its existence or to understand how this lubrication layer emerges and depends on ingredients and process parameters. Various experimental methods have been developed to analyze liquid phase migration in extrusion processes at laboratory scale [11, 5, 31, 32, 33, 34, 35]. These studies revealed that flow behaviour during extrusion is highly dependent on ingredient and process variables, including solid-liquid mixing ratio, particle size, deformability of particles, flow rate, pressure, and viscosity [15, 5, 36], yet stopped short of directly showing the existence of such a lubrication layer, let alone quantify it. Our surface liquid evaporation measurements [37], directly reveal (1) that this lubrication layer exists, (2) that it is only several micrometers thick, and (3) that the thickness is set by process parameters. We (4) confirm that lubrication tuning results in frictional changes by observing energy savings and temperature modulation. We demonstrate our results via pilot-scale pelleting plant trials, with realistic ingredient streams as typically used in biofuel and animal feed production. Our results can be translated to similar unit operations used in ceramics, foods, and pharmaceuticals production.
Experimental approach
Pellets are formed in a unit operation known as the pellet extrusion process (Fig. 1a), where particle mixtures, often from multiple ingredients, are pressed by a roller through cylindrically shaped holes in a metal ring-die — see Fig. 2a. The pellet extrusion process transforms the loose and low-density particle mixture into densely packed solid pellets.
@@
Unlike hard inorganic particle pastes used in e.g., concrete extrusion, organic plant-based particles used in food and feedstuff production exhibit hygroscopic properties, are capable of binding or releasing water, and undergo deformation and physico-chemical transitions while being processed. Consequently, the interaction mechanism and physical forces at the boundary layer between the metal extrusion die and the food/feed paste [2] are expected to differ from those of hard particles which for example are subject to excluded volume effects at the die-wall [7, 5, 8, 38].
Paste generation
@@
Prior to pelletization, conditioning uniformly mixes the ingredient particles and steam. The steam condenses upon cooling, transferring its latent heat into the particle mix and forms water [1, 39, 40]. Consequently, the steam conditioning process sets the temperature and amount of water present in the paste before it enters the extrusion die. It is known that increasing the water content in the mixture lowers the mechanical energy required during pellet extrusion [40]. While the water-induced lubrication in pellet extrusion is evident, the underlying mechanism remains unclear. The observed layer is too thin to be related to a particle migration phenomena, although such effects can also significantly change effective paste viscosity [41, 42]. Hitherto, one main assumption is that the paste during pellet extrusion internally retains its uniform moisture distribution $\phi_{\rm liquid}$ (see Fig. 1b) as it had right after the mixing/conditioning step [43, 44]. We show that this is not the case: the lubricating effects of added water is caused by the formation of a thin lubricating film that partially fills the pores between the metal die and the sliding pellet (Fig. 2f), reducing the real contact area between those respective surfaces, as expected in the mixed-lubrication regime [45]. Our perspective also explains the observed reduction in equipment wear with added water [40].
$\lambda$: Lubrication layer quantified
@@
We provide evidence for the existence of the liquid surface layer through a pilot-scale pelleting trial. Approximately 18 tons of pellets were produced while systematically adjusting the conditioning settings, as described in the Appendix B. The fresh pellets were sampled immediately after extrusion and placed in a custom designed sample container permitting drying of the product (Fig. 2a). The drying method involves exposing a sample container of pellets to a continuous, known flux of fresh air and recording the inlet and outlet air temperature and relative humidity (RH), along with the sample mass [37]. In 14 out of 15 drying experiments, a saturation period was observed ($t_{\mathrm{sat}}$), during which the passing air reached saturation (Fig. 2b). Over this period, up to 40% of the water introduced by conditioning was lost through evaporation (Fig. 2c), corresponding to the moisture “flash-off” commonly observed in pelleting factories [1].
Evaporation of the surface layer thus causes the initial saturation of drying air (indicated in Fig. 2b). The amount of water evaporating during this period can be quantified as we know the shape and thus total surface area of the pellets in the drying experiment, and the drying mass and density. We so obtain the average water layer thickness, $\lambda$, on the pellets surfaces (Fig. 2d). For the full calculation, see the appendix. The thickness of the lubrication layer increases from a few micrometers, similar to the height of surface asperities on the metal die (Fig. 2e) in which case we expect frictional interactions that are also known as Stribeck boundary lubrication [27]. As we increase water content of the paste ($\Delta\mathrm{mc}$) by increasing the steam flow in the conditioning process, the layer thickness increases to several tens of micrometers. Compared to the smallest individual particles produced by the milling procedure, this layer remains very thin [10] (Fig. 2d).
Lubrication affects extrudate surface structure
From the thickness measurements, we understand that the water layer at the surface between the plug and die wall only partially separates the pellet from the die. Increasing the water transfer during steam conditioning reduces, the area of direct contact between the pellet and die surface, leading to less damage to the particles embedded in the pellet surface. We see evidence for this in Fig. 2e&f and to less wear in the die. More evidence is presented in the Appendix: A.4, Fig. A.6. As the lubrication layer grows, we so expect to enter the “mixed regime” lubrication in the Stribeck perspective, and a concomitant reduction in friction.
The presence of the lubrication layer also induces plug flow in the compressed biomass as there is no visible internal deformation, uniform density and negligible porosity (see supportive X-ray tomography evidence described in Appendix: A.3).
@@
These collective findings suggest that the lubrication layer in organic particulate materials may not arise from factors like particle size distribution and the excluded volume effect, as observed in hard particle pastes. Instead, it appears to be influenced by the roughness of the die wall. In this scenario, water flows into the die’s surface cavities and reduces the shearing of the particles at the paste-die-wall interface by occupying part of the cavity volume, thus making it inaccessible for the feed particles. Consequently, the real contact area between the paste and metal die reduces, and the energy use decreases, as shearing of the water layer is much more favorable than shearing of the feedstuff material along the rough metal wall. This hypothesis aligns with industrial findings indicating that new pellet dies require more energy during pellet extrusion than used dies [46]. We attribute this energy use reduction over time to a decrease in the paste-die-wall contact area as the die’s largest asperities smooth out through wear. By smoothing of the surface, the water film can cover a larger area (the ratio of $\lambda$ over the wall roughness increases), further reducing the real contact area between the paste and die [28].
@@
The ability to measure and monitor the lubrication layer thickness enables us to decouple process dependent parameters such as steam use, die-hole dimensions and production rate (the latter, together with die-geometry specifications, determines residence time) from ingredient mixture dependent parameters, such as its chemical composition, particle size, particle size distribution and thermomechanical properties [47]. We identify two important consequences: we can now identify the source of mechanical energy use of pellet extrusion in the friction at the paste-die-wall interface, and consequently reduce it. Second, by tuning the lubrication layer, we can adjust the amount of heat added during extrusion, giving control over product treatment and quality.
Figure 3: (a) During pellet extrusion the specific mechanical energy usage is exponentially proportional to the coefficient of friction and compression ratio of the pelleting die. (b) Increasing the steam use during steam conditioning from a low moisture content (0.0350.0350.035 kg steam per kg ingredients) to a higher moisture content (0.0530.0530.053 kg steam per kg ingredients), results in a lower energy usage due to a decrease in the amount of friction. This is consistent with previous observations [40, 1].